OpenStudy (alexh107):

Find an equation in standard form for the ellipse with the vertical major axis of length 6 and minor axis of length 4.

OpenStudy (alexh107):

For a vertical one I have the equation as \[(x-h)^2/b^2 + (y-k)^2 / a^2 = 1\]

OpenStudy (anonymous):

what seems to be the problem ?

OpenStudy (alexh107):

I'm not sure how to apply the equation to the information the problem gives.

OpenStudy (anonymous):

here b = 4/2 =2 and a=6/2 =3 apply to the equation... not sure what could be values for h and k from the question

OpenStudy (faiqraees):

h,k are the coordinates of the centre of the ellipse

OpenStudy (alexh107):

Is that 0?

OpenStudy (faiqraees):

yes since its an standard equation

OpenStudy (alexh107):

Is it x^2/4 + y^2/9 = 1

OpenStudy (alexh107):

@FaiqRaees

OpenStudy (faiqraees):

yeah

OpenStudy (alexh107):

Okay, thank you!